Suppose you decide to travel to a star 65 light-years away with a speed that tells you the distance is only 25 light-years. How many years would it take to make the trip?


1
Expert's answer
2022-03-28T16:13:36-0400

"t=t_0\sqrt{1-\frac{v^2}{c^2}},"

"v=c\sqrt{1-\frac{t^2}{t_0^2}}=0.923c,"

"T=\frac sv=70.4" years.